Demonstrate the elimination method for a different system of equation : y=2x+4 and y=-3x-1

Answer by davethejackal(28)

y= 2x+4 [1]
y=-3x-1 [2]

Subtract [1] from [2]

Left hand side of [1] - left hand side of [2] = Right hand side of [1] - right hand side of [2]

Hence y-y = 2x - (-3x) +4 - (-1)
Implies 0 = 5x +5
Rearranging x = 5/(-5) = -1
Substitute x=-1 in [1], gives y = 2(-1)+4 = 2
Checking, substitute x=-1 in [2], gives y = -3(-1)-1 = 2
hence x = -1, y = 2